OpenStudy (anonymous):

The height of a parallelogram is three times it's base. If the area of the parallelogram is 972 cm squared, find the base and the height.

OpenStudy (anonymous):

Options: base=27cm, height =36cm b=324 cm, h=3 cm b=18 cm, h= 54 cm b=54 cm, h=162 cm

OpenStudy (campbell_st):

ok... this is what you know when you draw the picture  so then the area of a parallelogram is A = base x height \[972 = b \times 3b\] or \[972 = 3b^2\] just solve for b

OpenStudy (anonymous):

972=3b^2 /3 /3 324=b^2 18=b ?

OpenStudy (campbell_st):

that seems to be correct.. so the base is 18 and height = 3x18

OpenStudy (anonymous):

Ok thanks!